$75 is charged as a simple interest on a loan which is taken for 3 years at an interest rate of 5% per annum. What was the amount of loan in dollar?
Answer by Theo(13342)

simple interest loan formula is:
i = p * r * t
f = p + i
i = interest
p = principal
f = future value
you are given that i = 75, t = 3 years, r = .05 per year.
i = p * r * t becomes 75 = p * .05 * 3
solve for p to get p = 75 / (3 * .05) = 500
i = p * r * t becomes i = 500 * .05 * 3 which becomes i = 75, confirming the value of p is correct.
your solution is that the amount of the loan is 500.